OpenStudy (anonymous):

Write the equation of a circle whose center is at the origin and contains the point (7, 0).

OpenStudy (anonymous):

\[x^2+y^2=49\]

OpenStudy (anonymous):

from the picture we can see that the radius of the circle is 7, therefore the equation of the circle is\[x^2 + y^2 = 7^2\]\[x^2 + y^2 = 49\]